Help please!
What is the total energy, in MeV, of
A proton traveling at 98.3 % of the speed of light? VOI ΑΣφ ? E- Ε ΤΑ Mer
An electron traveling at 98.3 % of the speed of light? ΜΗ ΑΣΦ MeV
Answer #1

lev= Joule = 6.242X10 v=98-3 c -27 100 mp=1.67X10 kg v= 0.9830 Ep = 1 mp v2 Ep= I x1-67X1024x(@9830) 2 2 -27416 Ep=1_XE64x10Energy in. Men I Mev - lorer t Ep = 45.327X10 106 Aon EZESH = d7 Are Part Ime = 9.1X10-31 = 0.9836 7 KE = 2 31+16 KE- I me v2 I mex(0.98302 { x 9-141631x(0.9833 (3x103)2 KE= 39-5695 X10 KE-39.5695810l